An electric train starting from rest attains a maximum
speed of 200 kmph in 20 seconds. Determine :
(i) Its acceleration assuming it to be uniform
(ii) Distance covered during this time period
(iii) Its velocity 15 seconds after staring from rest
1
Expert's answer
2019-12-16T10:55:09-0500

From the formula for univariate movement

"v= v_0+a \cdot t"

where by the condition of the problem "v= 200 {km\/h}=55.56{m\/s}" "v_0=0"

1)define acceleration

"a=\frac{v-v_0}{t}=\frac{55.56-0}{20}=2.778[m/{s^2}]"

2) Distance traveled is

"s= \frac{a \cdot t^2}{2}=\frac{2.778 \cdot 20^2}{2}=555.6m"

3)Its speed after 15 seconds from the start of movement

"v=a \cdot t_1=2.778 \cdot 15=41.67{m/s}"